# Making Language Models Better Tool Learners with Execution Feedback

### 10 Feb 2024 v2

@@ Abstract
Tools serve as pivotal interfaces that enable humans to understand and reshape the environment. With the advent of foundation models, AI systems can utilize tools to expand their capabilities and interact with the real world. Existing tool learning methodologies, encompassing supervised fine-tuning and prompt engineering approaches, often induce large language models to utilize tools indiscriminately, as complex tasks often exceed their own competencies. However, introducing tools for simple tasks, which the models themselves can readily resolve, can inadvertently propagate errors rather than enhance performance. This leads to the research question: can we teach language models when and how to use tools? To meet this need, we propose Tool leaRning wIth exeCution fEedback (Trice), a two-stage end-to-end framework that enables the model to continually learn through feedback derived from tool execution, thereby learning when and how to use tools effectively. Experimental results, backed by further analysis, show that Trice can make the large language model selectively use tools by improving the accuracy of tool usage while enhancing insufficient tool learning and mitigating excessive reliance on tools [^1] Code and datasets are in https://github.com/zjunlp/TRICE..
@@
Shuofei ${}^{\spadesuit}$, Honghao ${}^{\spadesuit}$, Chengfei ${}^{\clubsuit}$, Qianghuai ${}^{\clubsuit}$, Huajun ${}^{\spadesuit\heartsuit}$, Ningyu ${}^{\spadesuit}$thanks:     Corresponding Author. ${}^{\spadesuit}$ Zhejiang University ${}^{\heartsuit}$ Donghai Laboratory ${}^{\clubsuit}$ Alibaba Group {shuofei,guihonghao,huajunsir,zhangningyu}@zju.edu.cn
@@ 3 Methodology
The data preparation follows the principles outlined in Eq.1&2. Given a raw initial dataset $\mathcal{D}_{\textrm{init}}=\{(q,a)\}_{i=1}^{|\mathcal{D}_{\textrm{init}}|}$ from the benchmark, we utilize LLM without fine-tuning to generate predictions. Since we do not have gold labels for tool APIs, we employ ChatGPT (OpenAI, 2022) to generate pseudo-labels under few-shot prompting. Specifically, we generate tool API labels $t=\mathtt{tool_{name}(tool_{input})}$ for questions where the generated predictions are incorrect. For questions with correct predictions, we directly set $t=\mathtt{None}$ to indicate that tool APIs are not required. We design particular instructions $s$ tailored to each task. In the end, we obtain $\mathcal{D}_{\textrm{tool}}=\{(s,q,t,a)\}_{i=1}^{|\mathcal{D}_{\textrm{tool}}|}$ according with Eq.1&2 containing the tool demand information of the specific LLM as we desire [^2] For more details of data preparation, please refer to A.2..
@@ 4.1 Experimental Settings
Tasks.
As shown in Table 2, we mainly evaluate our method on four tasks with each task specified to an external tool. Due to limited computational resources, we randomly sample train and test sets from each dataset to reduce the data scale. We display the detailed data distribution for each task in Figure 11. Following Schick et al. (2023), we use a more lenient evaluation criterion than exact match. We simply check for the last number predicted by the model for the math reasoning task and check whether the correct answer is within the first twenty words for other tasks.
@@
We collect five responses for each question from four different models, e.g. ChatGPT, InstuctGPT, Vicuna-7B, Alpaca-7B, and the output regulated in $\mathcal{D}_{\textrm{tool}}$ as the gold response. To differentiate whether or not to use tools among candidate responses, we compel ChatGPT and InstructGPT to utilize tools while allowing Alpaca and Vicuna to make the choice of using tools. For ChatGPT and InstructGPT, we prompt them with instructions and few-shot examples, and for Alpaca-7B and Vicuna-7B, we fine-tune them on $\mathcal{D}_{\textrm{tool}}$ with LoRA (Hu et al., 2022) for a few steps in order to equip them with initial abilities for question answering and tool generation [^3] For more details of candidate response generation, please refer to A.3..
@@
All the models are trained for 5 epochs in stage-I and 2 epochs in stage-II. We use the learning rates of {2e-5, 1e-4, 3e-4} for ChatGLM-6B and {2e-5, 1e-4} for Alpaca-7B and Vicuna-7B. The $\alpha$ is set to {0.01, 0.1, 1} for all the models. The detailed hyper-parameters we use are shown in A.4. Since sampling responses and training are separated, our whole training procedure only needs to load one model. All our training can be completed on one 80G A800 GPU within 10 hours.
@@ 4.2 Main Results
Within each task, we train the model to learn the corresponding tool as shown in Table 2, thereby evaluating the model’s proficiency in handling a single tool. From the rows labeled Trice-split in Table 3, it is evident that training by Trice, Alpaca and Vicuna perform on par with GPT-3.5, exhibiting only a slight decrease of $\downarrow$1.3and $\downarrow$0.4on average. Meanwhile, across all backbone models, Trice-split demonstrates significant improvements compared to the prompt-based baselines, surpassing the few-shot setting with $\uparrow$14.0for ChatGLM, $\uparrow$15.3for Alpaca, and $\uparrow$11.9for Vicuna. This indicates that Trice consistently empowers LLMs to use tools effectively, irrespective of the model architecture and scale. Moreover, whether it is completely independent (0% Tool) or dependent (100% Tool) on tools, supervised fine-tuning fails to beat Trice-based training, which highlights the necessity and efficacy of judicious tool learning.
@@
Table 4: Performance to unseen datasets and tools.
@@ 5 Discussion
One particularly challenging issue in tool learning is the problem of knowledge conflicts (Qin et al., 2023a) which may derive from the conflicts between model knowledge and augmented knowledge from tools. LLMs need to have the ability to differentiate knowledge from various sources and discern which ones are valuable, which ones are irrelevant, and even which ones may be harmful. Our approach leverages the feedback loop of trial and error (see Figure 2) to learn when to use tools and when not to. The model learns to recognize situations where relying solely on its intrinsic knowledge may not be sufficient and utilizing tools is more reliable. Similarly, it learns to identify scenarios where its own learned knowledge is capable of solving the problem without the need for extensive tool usage (see Figure 4).
@@
Recent NLP witnesses a rapid advancement in interactive learning (Wang et al., 2023). Collaboration among multi-agents (Lin et al., 2023; Liang et al., 2023) and learning from feedback (Chen et al., 2022; Ichter et al., 2022) are the keys to achieving general embodied intelligence as of now. Our approach is a preliminary endeavor to explore the incorporation of embodied methods into tool learning. By leveraging feedback obtained through interactions between the environment (tools) and multi-agents with varying capabilities, we enable language models to learn more desirable execution strategies (see Figure 5).

### 14 Mar 2024 v3

@@ Abstract
Tools serve as pivotal interfaces that enable humans to understand and reshape the environment. With the advent of foundation models, AI systems can utilize tools to expand their capabilities and interact with the real world. Existing tool learning methodologies, encompassing supervised fine-tuning and prompt engineering approaches, often induce large language models to utilize tools indiscriminately, as complex tasks often exceed their own competencies. However, introducing tools for simple tasks, which the models themselves can readily resolve, can inadvertently propagate errors rather than enhance performance. This leads to the research question: can we teach language models when and how to use tools? To meet this need, we propose Tool leaRning wIth exeCution fEedback (Trice), a two-stage end-to-end framework that enables the model to continually learn through feedback derived from tool execution, thereby learning when and how to use tools effectively. Experimental results, backed by further analysis, show that Trice can make the large language model selectively use tools by improving the accuracy of tool usage while enhancing insufficient tool learning and mitigating excessive reliance on tools [^1] Code: https://github.com/zjunlp/TRICE..
@@
Shuofei Qiao1,2, Honghao Gui1,2, Chengfei Lv4, Qianghuai Jia4, Huajun Chen1,2,3, Ningyu Zhang1,2thanks:     Corresponding Author. 1College of Computer Science and Technology, Zhejiang University 2ZJU-Ant Group Joint Research Center for Knowledge Graphs, Zhejiang University 3ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University 4Alibaba Group {shuofei,guihonghao,huajunsir,zhangningyu}@zju.edu.cn
@@ 3 Methodology
The data preparation follows the principles outlined in Eq.1&2. Given a raw initial dataset $\mathcal{D}_{\textrm{init}}=\{(q,a)\}_{i=1}^{|\mathcal{D}_{\textrm{init}}|}$ from the benchmark, we utilize LLM without fine-tuning to generate predictions. Since we do not have gold labels for tool APIs, we employ ChatGPT (OpenAI, 2022) to generate pseudo-labels under few-shot prompting. Specifically, we generate tool API labels $t=\mathtt{tool_{name}(tool_{input})}$ for questions where the generated predictions are incorrect. For questions with correct predictions, we directly set $t=\mathtt{None}$ to indicate that tool APIs are not required. We design particular instructions $s$ tailored to each task. In the end, we obtain $\mathcal{D}_{\textrm{tool}}=\{(s,q,t,a)\}_{i=1}^{|\mathcal{D}_{\textrm{tool}}|}$ according with Eq.1&2 containing the tool demand information of the specific LLM as we desire [^2] For more details of data preparation, please refer to Appendix A.2..
@@ 4.1 Experimental Settings
Tasks and Tools.
As shown in Table 2, we mainly evaluate our method on four tasks with each task specified to an external tool. Due to limited computational resources, we randomly sample train and test sets from each dataset to reduce the data scale. We display the detailed data distribution for each task in Figure 11. Following Schick et al. (2023), we use a more lenient evaluation criterion than exact match. We simply check for the last number predicted by the model for the math reasoning task and check whether the correct answer is within the first twenty words for other tasks. The QA model we use for LAMA is a retrieval-augmented LM fine-tuned on Natural Questions (Kwiatkowski et al., 2019) named Atlas (Izacard et al., 2023). We use the 600M parameter NLLB (Costa-jussà et al., 2022) as our machine Translation model for Multilingual QA.
@@
We collect five responses for each question from four different models, e.g. ChatGPT, InstuctGPT, Vicuna-7B, Alpaca-7B, and the output regulated in $\mathcal{D}_{\textrm{tool}}$ as the gold response. To differentiate whether or not to use tools among candidate responses, we compel ChatGPT and InstructGPT to utilize tools while allowing Alpaca and Vicuna to make the choice of using tools. For ChatGPT and InstructGPT, we prompt them with instructions and few-shot examples, and for Alpaca-7B and Vicuna-7B, we fine-tune them on $\mathcal{D}_{\textrm{tool}}$ with LoRA (Hu et al., 2022) for a few steps in order to equip them with initial abilities for question answering and tool generation [^3] For more details of candidate response generation, please refer to Appendix A.3..
@@
We fine-tune all our models with LoRA Hu et al. (2022) in the format proposed in Alpaca . All the models are trained for 5 epochs in stage-I and 2 epochs in stage-II. We use the learning rates of {2e-5, 1e-4, 3e-4} for ChatGLM-6B and {2e-5, 1e-4} for Alpaca-7B and Vicuna-7B. The $\alpha$ is set to {0.01, 0.1, 1} for all the models. The detailed hyper-parameters we use are shown in Appendix A.4. Since sampling responses and training are separated, our whole training procedure only needs to load one model. All our training can be completed on one 80G A800 GPU within 10 hours.
@@ 4.2 Main Results
Within each task, we train the model to learn the corresponding tool as shown in Table 2, thereby evaluating the model’s proficiency in handling a single tool. From the rows labeled Trice-split in Table 3, it is evident that training by Trice, Alpaca and Vicuna perform on par with GPT-3.5, exhibiting only a slight decrease of $\downarrow$1.3and $\downarrow$0.4on average. Meanwhile, across all backbone models, Trice-split demonstrates significant improvements compared to the prompt-based baselines, surpassing the few-shot setting with $\uparrow$14.0for ChatGLM, $\uparrow$15.3for Alpaca, and $\uparrow$11.9for Vicuna. This indicates that Trice consistently empowers LLMs to use tools effectively, irrespective of the model architecture and scale (ChatGLM-6B is encoder-decoder, while Alpaca-7B and Vicuna-7B are decoder-only). Moreover, whether it is completely independent (0% Tool) or dependent (100% Tool) on tools, supervised fine-tuning fails to beat Trice-based training, which highlights the necessity and efficacy of judicious tool learning.
@@
Table 4: Performance to unseen datasets and tools. The presence of an empty value in the Unseen Tool TRICE-APLIT section is due to our reliance on the generalization achieved through mixed-tool training (MIX). Testing the new tool individually through separate training (SPLIT) using any single tool would not be suitable.
@@ 5 Discussion
In tool learning, LLMs manipulate tools and respond to users conditioned on a variety of knowledge sources. One particularly challenging issue in tool learning is the problem of knowledge conflicts (Qin et al., 2023a) which may derive from the conflicts between model knowledge and augmented knowledge from tools, and among augmented knowledge from different tools. This may lead to a lack of explainability in model prediction and planning. LLMs need to have the ability to differentiate knowledge from various sources and discern which ones are valuable, which ones are irrelevant, and even which ones may be harmful. This ability becomes even more critical in highly specialized fields such as biomedical research and legal assistance, where the accuracy and reliability of knowledge are of utmost importance. Our approach leverages the feedback loop of trial and error (see Figure 2) to learn when to use tools and when not to. The model learns to recognize situations where relying solely on its intrinsic knowledge may not be sufficient and utilizing tools is more reliable. Similarly, it learns to identify scenarios where its own learned knowledge is capable of solving the problem without the need for extensive tool usage. This learning process allows the model to adapt and make informed decisions about when to rely on its own capabilities and when to utilize tools effectively (see Figure 4).
@@
Recent NLP has witnessed rapid advancement in interactive learning which considers language models as agents capable of observing, acting, and receiving feedback in a loop with external objects such as humans, knowledge bases, tools, models, and environments (Wang et al., 2023). Collaboration among multi-agents (Lin et al., 2023; Liang et al., 2023) and learning from feedback (Chen et al., 2022; Ichter et al., 2022) are the keys to achieving general embodied intelligence as of now. Our approach is a preliminary endeavor to explore the incorporation of embodied methods into tool learning. By leveraging feedback obtained through interactions between the environment (tools) and multi-agents with varying capabilities, we enable language models to learn more desirable execution strategies (see Figure 5). However, our current method is unable to learn the usage of multi-tool compositions. In the future, more sophisticated trial-and-error processes and feedback mechanisms will be necessary for LLMs to better utilize tools (e.g. learning multi-tool compositions) and even create new tools.
@@ 6 Conclusion
Acknowledgements
We would like to express gratitude to the anonymous reviewers for their kind comments. This work was supported by the National Natural Science Foundation of China (No. 62206246), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), CAAI-Huawei MindSpore Open Fund, Ningbo Natural Science Foundation (2021J190), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Tencent Rhino-Bird Open Research Fund, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.